# A Three-Level Whole-Body Disturbance Rejection Control Framework for Dynamic Motions in Legged Robots

### v2

@@ IX Conclusion
This paper presents an advanced control framework aimed at enhancing the stability and robustness of legged robots in the presence of various uncertainties, such as model uncertainties, external disturbances, and faults. The proposed framework integrates a MH-ESO to effectively estimate uncertainties and mitigate noise, ensuring accurate disturbance compensation. The introduction of the T-WB-DRC significantly improves robot performance by considering both uncertainty-free and uncertainty-affected dynamics, offering better payload transportation, disturbance rejection, and fault tolerance. Simulations on humanoid and quadruped robots demonstrate the framework’s versatility, while experimental trials on a quadruped robot validate its robustness and stability under a wide range of disturbance conditions. Overall, the proposed T-WB-DRC offers a promising approach to improving the performance of legged robots in dynamic and unpredictable environments.

### v3

@@ IX Conclusion
This paper presents an advanced control framework aimed at enhancing the stability and robustness of legged robots in the presence of various uncertainties, such as model uncertainties, external disturbances, and faults. The proposed framework integrates an MH-ESO to effectively estimate uncertainties and mitigate noise, ensuring accurate disturbance compensation. The introduction of the T-WB-DRC significantly improves robot performance by considering both uncertainty-free and uncertainty-affected dynamics, offering better payload transportation, disturbance rejection, and fault tolerance. Simulations on humanoid and quadruped robots demonstrate the framework’s versatility, while experimental trials on a quadruped robot validate its robustness and stability under a wide range of disturbance conditions. Overall, the proposed T-WB-DRC offers a promising approach to improving the performance of legged robots in dynamic and unpredictable environments.